If the area of the rectangle is 28 square feet, and the base is 7 feet. What is the height?
Found 2 solutions by negative49, psbhowmick:
Answer by negative49(13)

area of triangle = (1/2)(base)(height)
28 ft² = (1/2)(7 ft)(height)
28 ft² = (7/2 ft)(height)
28 ft² / 7/2 ft = height
8 ft = height

Answer by psbhowmick(878)

Height x Base = Area
Height x 7 = 28
Height = 28/7 = 4

So height of the rectangle is 4 ft.